OpenStudy (anonymous):

Simplify: 9 sq root 40 = 18 sq root 10????

OpenStudy (raden):

9 sqrt(40) = 9 * sqrt(4*10) = 9 * 2sqrt(10) = 18 sqrt(10) yeah, u are right

OpenStudy (anonymous):

thxs~~~~!

OpenStudy (raden):

welcome :*